2y - 3x = 0 , 2x - 3y = -10

Answer by checkley71(8403)

multiply the first equation by 2 & multiply the second equation by 3 & add them.
4y-6x=0
-9y+6x=-30
------------------
-5y=-30
y=-30/-5
y=6 answer.
2x-3*6=-10
2x-18=-10
2x=-10+18
2x=8
x=8/2
x=4 answer.
proof
2*6-3*4=0
12-12=0
0=0